How do I find the slope of any line perpendicular to 3x-4y=8?
Answer by nerdybill(7384)

First, determine slope of
3x-4y=8
by reorganizing into "slope-intercept" form:
3x-4y=8
-4y= -3x+8
y= (3/4)x-2
.
So, now we know the slope is 3/4
A line that is "perpendicular" to this must have a slope that is the "negative reciprocal" of the above:
Let m = new slope
then
(3/4)m = -1
3m = -4
m = -4/3
.
Solution: -4/3